# Why do Learning Rates Transfer? Reconciling Optimization and Scaling Limits for Deep Learning

### v1

@@ Appendix C Analysis of a Two-Layer Linear Network
$L(E,V)=\frac{1}{2p}\left\|\frac{1}{\gamma\sqrt{ND}}XEV-Y\right\|^{2}$ (7)
@@
C.1 Relationship between sharpness and residual, and proof of Lemma 4.1
Thus the Hessian blocks become:
@@
$\partial_{EE}$ $=\frac{1}{\gamma^{2}ND}E^{\top}E\in\mathbb{R}^{N\times N}$ (11)
@@
$\gamma^{2}\nabla^{2}\mathcal{L}(E,V)$ $=\underbrace{\frac{1}{ND}\begin{pmatrix}I_{D}\otimes VV^{\top}&E\otimes V\\ E^{\top}\otimes V^{\top}&E^{\top}E\end{pmatrix}}_{\mathcal{G}(E,V)}+\underbrace{\sqrt{\frac{\gamma}{ND}}\begin{pmatrix}0&I_{N}\otimes(w-w_{*})\\ I_{N}\otimes(w-w_{*})^{\top}&0\end{pmatrix}}_{\mathcal{R}(E,V)},$ (21)
@@
$\lambda_{\max}[\mathcal{R}(E,V)]=-\lambda_{\min}[\mathcal{R}(E,V)]=\sqrt{\frac{\gamma}{ND}}\|w-w^{*}\|.$
@@
$\lambda_{\max}[\Theta(E,V)]-\sqrt{\frac{\gamma}{ND}}\|w-w_{*}\|_{2}\leq\lambda_{\max}[\gamma^{2}\nabla^{2}\mathcal{L}(E,V)]\leq\lambda_{\max}[\Theta(E,V)]+\sqrt{\frac{\gamma}{ND}}\|w-w_{*}\|_{2}.$ (25)

### v2

@@ Appendix C Analysis of a Two-Layer Linear Network
$L(E,V)=\frac{1}{2}\left\|\frac{1}{\gamma\sqrt{ND}}XEV-Y\right\|^{2}$ (6)
@@
C.1 Relationship between sharpness and residual, and Gauss-Newton
The following bound leverages a Gauss-Newton decomposition and leads analytical insights supporting Sec. 4.1. Proof of Lemma 5.2, which we restate here.
Lemma (GN bound).
$\Theta(E,V)=e+v\cdot I_{D}$
and
$|\lambda_{\max}[\gamma^{2}\nabla^{2}\mathcal{L}(E,V)]-\lambda_{\max}[\Theta(E,V)]|\leq\sqrt{\frac{\gamma^{2}}{ND}}\|w-w_{*}\|_{2}.$
The result above is actually more general: it indicates that the eigenvalues of the positive semidefinite portion of the Hessian $\mathcal{G}=K^{\top}K$ are fully [^5] Simple application of the SVD decomposition. characterized by the eigenvalues a smaller matrix $\Theta=KK^{\top}$. Furthermore, in our model $\Theta$ has a closed form that depends only on the variables $e,v$.
Proof.
The Hessian blocks become:
@@ C.1 Relationship between sharpness and residual, and Gauss-Newton
$\partial_{VV}$ $=\frac{1}{\gamma^{2}ND}E^{\top}E\in\mathbb{R}^{N\times N}$ (10)
@@
$\gamma^{2}\nabla^{2}\mathcal{L}(E,V)$ $=\underbrace{\frac{1}{ND}\begin{pmatrix}I_{D}\otimes VV^{\top}&E\otimes V\\ E^{\top}\otimes V^{\top}&E^{\top}E\end{pmatrix}}_{\mathcal{G}(E,V)}+\underbrace{\sqrt{\frac{\gamma^{2}}{ND}}\begin{pmatrix}0&I_{N}\otimes(w-w_{*})\\ I_{N}\otimes(w-w_{*})^{\top}&0\end{pmatrix}}_{\mathcal{R}(E,V)},$ (12)
@@
$\lambda_{\max}[\mathcal{R}(E,V)]=-\lambda_{\min}[\mathcal{R}(E,V)]=\sqrt{\frac{\gamma^{2}}{ND}}\|w-w^{*}\|.$
@@
$\lambda_{\max}[\Theta(E,V)]-\sqrt{\frac{\gamma^{2}}{ND}}\|w-w_{*}\|_{2}\leq\lambda_{\max}[\gamma^{2}\nabla^{2}\mathcal{L}(E,V)]\leq\lambda_{\max}[\Theta(E,V)]+\sqrt{\frac{\gamma^{2}}{ND}}\|w-w_{*}\|_{2}.$ (16)